OpenStudy (anonymous):

5x-13=3x+6 solve for x pleasee help

OpenStudy (anonymous):

x=9.5

OpenStudy (anonymous):

could you show your work please

OpenStudy (anonymous):

5x-13=3x+6 5x=3x=6+13 2x=19 x=9.5

OpenStudy (anonymous):

thanks